im having a hard time with this problem. Find the Domain of g(x) = [x4 -2x3 + 7] / [3x2 – 10x – 8] and f(x) = [x+1]1/2 / x. Can someone help me?
Answer by stanbon(75887)

Find the Domain of g(x) = [x^4 -2x^3 + 7] / [3x^2 – 10x – 8]
Need to find out when the denominator is zero:
3x^2-10x-8=0
3x^2-12x+2x-8=0
3x(x-4)+2(x-4)=0
(x-4)(3x+2)=0
x=4 or x=-2/3
Domain is all real numbers except x=4 and x=-2/3
======================
and f(x) = [x+1]^(1/2) / x.
x in the denominator cannot be zero
x+1 must be >=0
x>=-1
Domain is all Real Numbers greater than or equal to -1 excluding zero.
===========
Cheers,
Stan H.